It takes an older pump four times as long to drain a certain pool as it does a newer pump. Working together it takes the two pumps three hours to drain the pool . How long will it take the newer pump to drain the pool working alone?
Answer by Boreal(15235)

newer pump takes x hours
older pump takes 4x hours
in 1 hour they drain (1/x) and (1/4x) respectively.
(3/x)+(3/4x)=1 full pool drained.
multiply by 4x
12+3=4x
x=(15/4) hours or 3.75 hours for new pump
in 3 hours, it drains 3/3.75 of the pool or 80%
the other pump drains 20% in that time